F(t)=sint×u(t-2) laplace transform
1
Expert's answer
2020-04-13T12:41:05-0400

"F(t)=sint\times u(t-2)"

Let "(t-2)=z\implies dt=dz"

So "F(z)=sin(z+2)\times u(z)"

Laplace will be "F(s)"

As "F(s)=\int^{\infin}_{0}e^{-s(z+2)}sin(z+2)dz"

Using by parts integration,

"F(s)=sin(z+2)\frac{e^{-s(z+2)}}{-s}+\frac{1}{s}\int^{\infin}_{0}cos(z+2)e^{-s(z+2)}dz"

Apply by parts integration again, we get

"F(s)=\frac{e^{-2s}(s\times sin2+cos2)}{s^2+1}"